OpenStudy (anonymous):

The number of milligrams of a drug that remains in a patient’s system after t hours is given by the function A(t) = Ie^rt. Juan was given 500 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 6 hours

jimthompson5910 (jim_thompson5910):

"Juan was given 500 milligrams" which means the initial amount I is I = 500, so A(t) = Ie^rt becomes A(t) = 500e^rt

OpenStudy (anonymous):

i got 111.58 mg

jimthompson5910 (jim_thompson5910):

Now in 1 hour, 20% of 500 mg leaves, so you're left with 80% of 500 or 0.8*500 = 400 mg This means A(t) = 500e^rt becomes 400 = 500e^(r*1) What do you get when you solve for r?

jimthompson5910 (jim_thompson5910):

111.58 mg is a bit short of what I'm getting

OpenStudy (anonymous):

is 133.90mg closeR?

jimthompson5910 (jim_thompson5910):

yes it is, but still not what I'm getting

jimthompson5910 (jim_thompson5910):

what did you get for r?

OpenStudy (anonymous):

the answer is 150.60 mg! :)

jimthompson5910 (jim_thompson5910):

is that what the book says?

OpenStudy (anonymous):

there is no book.

jimthompson5910 (jim_thompson5910):

oh so that's the computer says?

jimthompson5910 (jim_thompson5910):

that's what*

OpenStudy (anonymous):

thats the highest choice.

jimthompson5910 (jim_thompson5910):

is 131.07 a choice?

OpenStudy (anonymous):

133.90 is the closet to that number

jimthompson5910 (jim_thompson5910):

hmm maybe something is missing because 131.071 is what I'm getting

OpenStudy (anonymous):

im not sure i just checked the problem

jimthompson5910 (jim_thompson5910):

Is 104.86 a choice?

OpenStudy (anonymous):

the choices are 150.60 mg 133.0mg, 111.58mg,99.20mg

jimthompson5910 (jim_thompson5910):

yeah something has to be missing, I'm not getting any of those

OpenStudy (anonymous):

no thats the problem. perfectly

OpenStudy (kropot72):

This is exponential decay. The equation should be\[A(t)=Ie ^{-rt}\] Put r = 0.2 and the answer is 150.6mg

jimthompson5910 (jim_thompson5910):

You mean r = -0.2, I guess it's that simple...

OpenStudy (kropot72):

@jim_thompson5910 The exponent is -rt. You could give the fractional rate a negative sign and say the original equation was correct I guess. :)

jimthompson5910 (jim_thompson5910):

oh, in this case, they say that the exponent is just rt...hmm this problem is just off